Circular No. 9014

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

R CORONAE BOREALIS M. Templeton, AAVSO, informs us that this variable is apparently at

or near its historic minimum; a number of observers have found this star to be fainter than visual mag 14.0 since early Nov. 2008, and both visual and instrumental measures are now indicating V ≥ 14.5. R CrB began its current fading episode around 2007 July 6 and faded from m_{v} ∼ 6.0 to > 12.0 by 2007 Aug. 12 (cf. CBET 1052); the star has continued to fade for the past 17 months. The duration of the current episode and its depth are similar to that observed during the previous extreme fading episode that began around June 1963 and continued (with only one brief interruption) until around Dec. 1965, when R CrB faded to visual mag about 14.5.

COMETS C/2003 Q1, Q6; C/2008 R13, AND C/2008 S4–S8 (SOHO)

Further to IAUC 9011, additional Kreutz sungrazing comets have been found on SOHO website images. C/2008 R13 (which was inadvertently tabulated on IAUC 9011; peak magnitude ∼ 7.5), C/2008 S4 (peak mag ∼ 6.5–7), and C/2008 S5 (peak mag ∼ 7) were small and stellar in appearance. C/2008 S6 was tiny and stellar in appearance (mag ∼ 7.5) in C3 images, but quite diffuse and very faint in C2 images. C/2008 S7 was small and stellar in appearance (mag ∼ 6.5–7) in C3 images, but very diffuse and slightly elongated in C2 images. C/2008 S8 was stellar in appearance (mag ∼ 6.5–7) in C3 images, but diffuse with a very faint, diffuse, short tail in C2 images.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 S4 Sept. 17.571 11^{h}23.^{m}1 + 0^{o}54′ C3  RM 2008-X25
C/2008 S5 19.696 11 31.2 + 0 18 C3 RM 2008-Y04
C/2008 S6  28.388 12 09.3 − 3 43 C3/2 MK 2008-Y04
C/2008 S7  29.096 12 03.6 − 3 46 C3/2 JR 2008-Y04
C/2008 S8  30.179 12 08.4 − 4 12 C3/2 MK 2008-Y04

K. Battams adds that C/2003 Q1 and C/2003 Q6 (cf. IAUC 8339) were both tiny and stellar in appearance (peaking at mag ∼ 8–8.5); improved astrometry for these presumed comets was published on MPEC 2008-Y06.

η CARINAE

Corrigendum. On IAUC 9011, line 12, for reached −3204 km/s read reached −320 km/s

2009 January 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green